OpenStudy (brenna):

what is the maximum number of turns in the graph of f(x)=-x^4+3x^2

imqwerty (imqwerty):

The maximum number of turns will be equal to the number of roots of the equation f'(x) =0

OpenStudy (giitter):

lol for once i see you answering a question

imqwerty (imqwerty):

=] she asked for halp so i came lol don't spam tho

OpenStudy (giitter):

I wont im here to solve dis questions >:D

OpenStudy (kevin):

do your best, both of you :D

OpenStudy (giitter):

Here is how you solve this answer- Let's solve for f. fx=−x4+3x2 Step 1: Divide both sides by x. fx x = −x4+3x2 x f=−x3+3x our answer is f=−x3+3x Do what you love, love you do and thanks for post this question for me to answer! Feel free to pm me if you have any questions/concerns about my answer! ~Glitter

OpenStudy (brenna):

thanks so much for the help!!!

imqwerty (imqwerty):

Yw